OpenStudy (anonymous):

simplify: (csc^2(u) -1)sin^2(u)

OpenStudy (anonymous):

\[\csc(x)=\frac{1}{\sin(x)}\] so \[\csc^2(u)\times \sin^2(u)=1\] is a start

OpenStudy (anonymous):

multiplying out gives \[1-\sin^2(u)\] and since \(\cos^2(x)+\sin^2(x)=1\) that is the same as \(\cos^2(u)\)